OpenStudy (anonymous):

find the derivative of 1-10t/(5t)^1/2 I have tried this problem like about ten times and I can not get to the right answer.

OpenStudy (anonymous):

\[1-\frac{10t}{\sqrt{5t}}\] Is that your question. Just want to be sure.

OpenStudy (anonymous):

no it's \[\frac{ 1-10t }{ (5t)^1/2 }\]

OpenStudy (anonymous):

Oh okay. Then use the product rule to get: \[-\frac{10t+1}{2\sqrt{5} (t^\frac{3}{2})}\]

OpenStudy (anonymous):

I'll try that thanks.